# Fast Bilateral Teleoperation and Imitation Learning Using Sensorless Force Control via Accurate Dynamics Model

### v1

@@ Abstract
![[Uncaptioned image]](extracted/6606566/figures/thumbnail.png)
@@ 2 Methodology
where is the reference torque other than compensation, denotes parameter-identified manipulator dynamics functions, and denotes values estimated by an observer, which will be explained later. Conventionally, it was difficult to perform bilateral control using strict dynamic models such as (2) and (3). We propose to conduct parameter identification using the regressor matrix based on OpenSYMORO [28], which is software to derive efficient robot dynamics parameters, and the least squares method. Here, we consider the coordinate transformation of the joint angler velocity of the leader and follower into differences and averages.
@@
For teleoperation, the desired dynamics for the joint angle difference are fixed, which means heavy weight, significant friction, and high stiffness. On the other hand, the desired dynamics for the average joint angle are light and have no spring or damper. The desired dynamics can be written as a spring-mass-damper system like impedance control [29] on transformed coordinates [16] as below:
@@
where , , . Finally, the control law for each manipulator can be as follows:
@@ 2.2 Velocity and External Force Observer
Assuming that external forces remain constant (), the dynamics of the manipulator can be described in the following state equation form.
Minimal-order observers are often used to minimize the phase delay of external force estimation in disturbance observers [14][33]. Although several methods have been proposed for constructing minimal-order observers, here we explain a specific, simplified method. First, assuming that velocity can be observed, we will construct a full-order observer using a state-space model with only velocity and external force.
@@ 3.1 Teleoperation Performance
vs. 4ch Bilateral Control using fixed inertia. We compared the results obtained using fixed inertial parameters to those obtained when inertial variations were considered. Although control systems with fixed inertia (fixed gain) are simple and easy to stabilize and they are often used in manipulator control, their performance depends on posture and is not constant. The results of the experiment showed that considering inertial variations significantly reduced position errors.
@@ 3.2 Imitation Learning
It should be noted that a control system using position and force command values was employed during data collection. Since the control system differs between data collection and autonomous operation, a decrease in success rate when force information is not used in the output of policy was expected. However, differences in control systems or dynamics between data collection and autonomous operation are common when using data collection methods other than teleoperation, and this is not uncommon in imitation learning. Additionally, since the dataset’s quality can significantly affect imitation learning results, it is preferable to compare using the same dataset. Considering these points, we conducted this comparison.

### v2

@@ Abstract
![[Uncaptioned image]](extracted/6622156/figures/thumbnail.png)
@@ 2 Methodology
where is the reference torque other than compensation, denotes parameter-identified manipulator dynamics functions, and denotes values estimated by an observer, which will be explained later. Conventionally, it was difficult to perform bilateral control using strict dynamic models such as (2) and (3). We propose to conduct parameter identification using the regressor matrix based on OpenSYMORO [28], which is software to derive efficient robot dynamics parameters, and the least squares method. Here, we consider the coordinate transformation of the joint angular velocity of the leader and follower into differences and averages.
@@
For teleoperation, the desired dynamics for the joint angle difference are fixed, which means heavy weight, significant friction, and high stiffness. On the other hand, the desired dynamics for the average joint angle are light and have no spring or damper. The desired dynamics can be written as a spring-mass-damper system like impedance control [29] on transformed coordinates [16][30] as below:
@@
where , , and . Finally, the control law for each manipulator can be as follows:
@@ 2.2 Velocity and External Force Observer
Assuming that the differential of the external forces remain zero (), the dynamics of the manipulator can be described in the following state equation form.
Minimal-order observers are often used to minimize the phase delay of external force estimation in disturbance observers [14][34][35]. Although several methods have been proposed for constructing minimal-order observers, here we explain a specific, simplified method. First, assuming that velocity can be observed, we will construct a full-order observer using a state-space model with only velocity and external force.
@@ 3.1 Teleoperation Performance
vs. 4ch Bilateral Control using fixed inertia. We compared the results obtained using fixed inertial parameters to those obtained when inertial variations were considered. Although control systems with fixed inertia (fixed gain) are simple and easy to stabilize, and they are often used in manipulator control, their performance depends on posture and is not constant. The results of the experiment showed that considering inertial variations significantly reduced position errors.
@@ 3.2 Imitation Learning
It is worth noting that a control system utilizing position and force command values was employed during data collection. Since the control system differs between data collection and autonomous operation, a decrease in success rate was expected when force information is not used in the policy output. However, differences in control systems or dynamics between data collection and autonomous operation are common when using data collection methods other than teleoperation, and this is not uncommon in imitation learning. Additionally, since the dataset’s quality can significantly affect imitation learning results, it is preferable to compare using the same dataset. Considering these points, we conducted this comparison.